OpenStudy (anonymous):

Simplify. (-5.12) - (-4.7) + (7.021)

OpenStudy (anonymous):

-2.799

OpenStudy (anonymous):

(-5.12) - (-4.7) + (7.021) = -5.12 +4.7+7.021 = -5.12 + 11.721 = 6.601

OpenStudy (anonymous):

negative and negative make a positive:\[-5.12+4.7+7.021\]now just add them all:\[-.042+7.021\]\[=6.601\]

OpenStudy (anonymous):

@yummydum I got the same result. u can try other sums by my method, u will get the same results